OpenStudy (anonymous):

An oak tree that is 220 years old produced acorns at two different rates. Until it reached maturity, it produced an average of 900 acorns a year. However, once the tree reached maturity, it produced an average of 2200 acorns a year. One particular tree, Angel Oak, has produced about 289,000 acorns in its lifetime. At what age did it reach maturity?

OpenStudy (anonymous):

would i add the year and month ?

OpenStudy (anonymous):

@Lily2913 @UsukiDoll

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Redcan

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

i just dont get this

OpenStudy (zzr0ck3r):

\(t_1+t_2=220\) where \(t_1\) is the time that the first tree was young, and \(t_2\) is the time it was mature. make sense?

OpenStudy (anonymous):

wich is 220 years ? for t1?

OpenStudy (zzr0ck3r):

so a tree grows for 220 years, part of that time it is young, we call that time \(t_1\) and the rest of the time it is old, we call that time \(t_2\). So \(t_1+t_2=220\)

OpenStudy (anonymous):

so divide it by 2 ???

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

I am just asking if you understand that

OpenStudy (anonymous):

im sorry i really dont get this

OpenStudy (zzr0ck3r):

that if we add the time it was young and time it was old we get the total time?

OpenStudy (zzr0ck3r):

does that make sense/.

OpenStudy (anonymous):

okay i get that

OpenStudy (zzr0ck3r):

so \(900t_1\) is the amount of trees it produces when it is young.

OpenStudy (zzr0ck3r):

like if it was two years old it gave 1800 trees.

OpenStudy (zzr0ck3r):

right?

OpenStudy (anonymous):

okay right

OpenStudy (zzr0ck3r):

and similarly we have \(2200t_2\) gives the amount of trees when it is old

OpenStudy (zzr0ck3r):

if we add these together we have the total amount of trees. \(900t_1+2200t_2=289000\)

OpenStudy (zzr0ck3r):

the amount of trees it produced when it was young plus the amount of trees it produced when it was old is 289000

OpenStudy (anonymous):

okay i got that, i just dont know find the year

OpenStudy (zzr0ck3r):

so we have two equations \(900t_1+2200t_2=289000\\t_1+t_2=220\) Solve for \(t_2\) in the second equation and you get \(t_2=200-t_1\) right? We are almost there.

OpenStudy (anonymous):

how did you get that ?

OpenStudy (zzr0ck3r):

\(t_1+t_2=220\) so we just subtract \(t_1\) from both sides and get \(t_2=220-t_1\)

OpenStudy (anonymous):

t1 which is 900 ?

OpenStudy (zzr0ck3r):

no

OpenStudy (anonymous):

can you maybe sketch it out

OpenStudy (zzr0ck3r):

you agreed that \(t_1+t_2=220\) and you also agreed \(900t_1 + 2200t_2=290000\) now all I did was take \(t_1+t_2=200\) and subtract \(t_1\) from both sides. So now I have \(900t_1+2200t_2=289000\\t_2=220-t_1\) Take the second equation and plug it in to the first one. \(900t_1+2200(200-t_1)=289000\) Now lolve for \(t_1\)

OpenStudy (zzr0ck3r):

solve*

OpenStudy (zzr0ck3r):

the very first line in that should say 289000 not 290000

OpenStudy (anonymous):

omg i really dont understand

OpenStudy (anonymous):

whats t1 ?

OpenStudy (anonymous):

so we have to figure out what t1 and t2 equals ?

OpenStudy (anonymous):

So in zzrocks3r's notation, t1 = 0-tm, and t2=220-tm.

OpenStudy (anonymous):

tm ?

OpenStudy (anonymous):

okay so whats t1 ?

OpenStudy (anonymous):

The time the tree is producing 900 acorns/year ---> t1

OpenStudy (anonymous):

is there a easier way to do this because im very very very confused

OpenStudy (anonymous):

so whats the equation ? and how do i solve it

OpenStudy (anonymous):

??

OpenStudy (anonymous):

ok ?